Consider a router that interconnects IPv4 three subnets: Subnet 1, Subnet 2, Subnet 3. Suppose all of the interfaces in each of these three subnets are required to have the prefix 172.12 /16. Also suppose that Subnet 1 is required to support at least 30000 interfaces, Subnet 2 is to support at least 2000 interfaces, and Subnet 3 is to support at least 10000 interfaces. Provide three network addresses (of the form a.b.c.d/x) that satisfy these constraints
Answer #1

The address given to us is 172.12.0.0 /16, which means we have 16 bits for the host id.

Further, when we have to divide this into 3 subnets with the given configuration, we will first arrange the subnets in descending order based on the number of interfaces required. So we will have

Subnet 1 with 30000 interfaces,

Subnet 3 with 10000 interfaces, and

Subnet 2 with 2000 interfaces,

So now, we have the order in which we will create the subnets. We will divide the given address into our 3 subnets in the following manner,

1. For subnet 1

The number of interfaces required = 30000

The number of bits required to allocate an address to these interfaces = 215.

Therefore, out of 16 host id bits, 15 bits will be allocated to Subnet 1 itself.

We will allocate the first bit of the third byte to (0)2 for this subnet. Consequently, we will get the IP addresses as

Initial IP address = 172.12.0.0 /17

Final IP address = 172.12.127.255 /17

2. For Subnet 3

The number of interfaces required = 10000

The number of bits required to allocate an address to these interfaces = 214.

Therefore, out of the remaining 15 host id bits, 14 bits will be allocated to Subnet 2.

We will allocate the first two bits of the third byte to (10)2 for this subnet. Consequently, we will get the IP addresses as

Initial IP address = 172.12.128.0 /18

Final IP address = 172.12.191.255 /18

3. For Subnet 2

The number of interfaces required = 2000

The number of bits required to allocate an address to these interfaces = 211.

Therefore, out of the remaining 14 host id bits, 11 bits will be allocated to Subnet 2.

We will allocate the first five bits of the third byte to (11000)2 for this subnet. Consequently, we will get the IP addresses as

Initial IP address = 172.12.192.0 /21

Final IP address = 172.12.199.255 /21

Therefore, the three network addresses are:-

172.12.0.0 /17 for Subnet 1

172.12.192.0 /21 for Subnet 2

172.12.128.0 /18 for Subnet 3